OpenStudy (anonymous):

What is the best method (system of equations) to find the solution of the following equation(s)? a. 3x - 9y = 3 6x - 3y = -24

OpenStudy (sasogeek):

elimination

OpenStudy (cwrw238):

elimination method is good here multiply first equation by -2 and add

OpenStudy (ash2326):

substitution would be easy divide first equation by 3 x-3y=1 x=3y+1 substitute this in second 6(3Y+1)-3y=-24 18y+6-3y=-24 15y=-30 y=-2 x=-2*3+1 x=-5 and y=-2

OpenStudy (anonymous):

Why is elimiation the best method? (Short answer please )

OpenStudy (sasogeek):

it's faster

OpenStudy (cwrw238):

yea - i agree with that

OpenStudy (sasogeek):

get rid of x or y, solve for the one that's left, plug your answer into the other equation, solve for the other variable, you're done